OpenStudy (mayaal):

find the period of y= -3cos x

OpenStudy (anonymous):

same as the period of cosine

OpenStudy (anonymous):

the \(-3\) out front has nothing to do with the period

OpenStudy (mayaal):

thanku for helping me again @satellite73 :)

OpenStudy (anonymous):

2pi

OpenStudy (anonymous):

yw

OpenStudy (mayaal):

how did u get the 2pi @henryarias5 ?

OpenStudy (anonymous):

because the is the period of cosine

OpenStudy (anonymous):

also the period of sine

OpenStudy (anonymous):

tangent is a different story, it has period \(\pi\)

OpenStudy (imstuck):

The -3 is indicative of the fact that the amplitude is 3 and it is reflected across the x axis.

OpenStudy (mayaal):

and is there a formula to find the period?

OpenStudy (mayaal):

anybody?

OpenStudy (anonymous):

Yes!! it is \(\dfrac{2\pi}{\text{the number stands in the front of x}}\) For example: sin x (the number in the front of x is 1) --> period = \(\dfrac{2\pi}{1}=2\pi\) sin (2x) (the number in the front of x is 2) --> period = \(\dfrac{2\pi}{2}=\pi\) cos (5x) (the number in the front of x is 5) --> period = \(\dfrac{2\pi}{5}\)

OpenStudy (mayaal):

thnks @ooops